My talk covered one of the main results in the paper: J.-L. Colliot-Thélène, J.-J. Sansuc et Sir Peter Swinnerton-Dyer) Intersections of two quadrics and

Châtelet surfaces, I, J. für die reine und angew. Math. (Crelle) 373 (1987) 37-107; II, ibid. 374 (1987) 72-168.

Other papers relevant to my talk at the Winter School are: J.-L. Colliot-Thélène et J.-J. Sansuc, La descente sur les variétés rationnelles. In Journées de géométrie

algébrique d’Angers (Juillet 1979),édité par A.Beauville, Sijthof and Noordhof (1980) 223-237.

J.-L. Colliot-Thélène et J.-J. Sansuc, La descente sur les variétés rationnelles II, Duke Math. J. 54 (1987) 375-492.

J.-L. Colliot-Thélène, Surfaces rationnelles fibrées en coniques de degré 4, Séminaire de théorie des nombres de Paris 88-89. Progress in Mathematics, Birkhäuser (1990), p. 43-55.

J.-L. Colliot-Thélène et P. Swinnerton-Dyer, Hasse principle and weak approximation for pencils of Severi-Brauer and similar varieties J. für die reine und ang. Mathematik (Crelle) 453 (1994) 49-112.

J.-L. Colliot-Thélène, A. N. Skorobogatov et Sir Peter Swinnerton-Dyer, Rational points and zero-cycles on fibred varieties : Schinzel’s hypothesis and Salberger’s device, Journal für die reine und angewandte Mathematik (Crelle) 495 (1998) 1-28.

J.-L. Colliot-Thélène, A. N. Skorobogatov et Sir Peter Swinnerton-Dyer, Hasse principle for pencils of curves of genus one whose Jacobians have rational 2-division points, Inventiones math. 134 (1998) 579-650.

J.-L. Colliot-Thélène and A. Skorobogatov, Descent on fibrations over P_{k}^{1} revisited, November 1998. To appear in Mathematical Proceedings of the Cambridge Philosophical Society (2000).

Surveys : J.-L. Colliot-Thélène, Arithmétique des variétés rationnelles et problèmes birationnels. Proceedings of

the International Congress of Mathematicians, Berkeley, California 1986, (1987), Tome I, 641-653.

J.-L. Colliot-Thélène, L’arithmétique des variétés rationnelles. Annales de la Faculté des sciences de Toulouse, vol. I, numéro 3 (1992) 295-336.

J.-L. Colliot-Thélène,The Hasse principle in a pencil of algebraic varieties, Proceedings of the Tiruchirapalli conference (India, January 1996), ed. M. Waldschmidt and K. Murty, Contemporary Mathematics, 210 (1998) 19-39.

Almost all these papers have appeared, as you may see. Here is the address of my web page : http://www.math.u-psud.fr/˜ colliot/ There people will find links to preprints.